# On the Feasibility of Using LLMs to Execute Multistage Network Attacks

### v2

@@ 1. Introduction
Dumbbell B Has the same topology and goal as Dumbbell A. Each web server has credentials to databases, but they require privileged access. In addition, the data on each database requires privileged access. Exfiltrate all critical data. 30
Table 1. We implement 10 multistage environments. The Equifax and Colonial pipeline environments are based on real attacks (equifax_report, ; colonial_pipeline_techtarget, ). The Enterprise environments are inspired by topologies of enterprise networks (ciscoEnterpriseNetwork, ; ibmTreeNetwork, ). The Chain, Star, and Dumbbell environments are in prior work (mirage, ; ringNetwork2013technique, ; ferguson2021_deception_psychology, ).
2. Related Work and Motivation
We start with a brief overview about related work in LLMs executing attacks. Then, we address a key blind spot in prior work—understanding how LLMs perform at executing multistage network attacks. To this end, we evaluate popular LLMs in 10 multistage attack environments ranging from 25–50 hosts.
2.1. Related work on LLMs executing attacks
Prior studies evaluate LLMs at solving CTF-style challenges (pentestgpt, ; cyberseceval3, ; o1systemcard, ; zhang2024cybench, ; anthropic_AISI, ; penheal, ; happe2023getting, ; yang2023language, ; shao2024nyu_ctf, ; intercode_ctf, ; anurin2024catastrophic, ; fang2024teams, ; shao2024empirical, ). Many of these CTFs are challenge problems related to security but do not involve infecting a host (e.g., finding an XSS vulnerability or solving a cryptography challenge (zhang2024cybench, ; pentestgpt, ; o1systemcard, )). At most, the challenges are single host attacks that involve infecting a single host (pentestgpt, ; cyberseceval3, ; o1systemcard, ; zhang2024cybench, ; anthropic_AISI, ). While some of these challenges require multiple low-level steps (e.g., identify a remote service, discover a vulnerability, then exploit the vulnerability) (zhang2024cybench, ; o1systemcard, ; pentestgpt, ), they do not involve multiple hosts and subnetworks. We refer to challenges that involve multiple hosts and subnetworks as multistage network attacks. Existing efforts have not tackled multistage attacks, which is the focus of our work.
At a high level, prior work on using LLMs in attack challenges fall in two classes: (1) fully autonomous (e.g., (cyberseceval3, ; o1systemcard, ; anthropic_AISI, ; zhang2024cybench, )) and (2) human-assisted LLM attack systems (e.g., (zhang2024cybench, ; pentestgpt, )). The autonomous frameworks instruct LLMs to attack the environment by outputting command line instructions. Then, a second program automatically extracts commands from the LLM’s response and executes it on a computer with access to the environment.
In contrast, human-assisted LLM-based attack tools such as PentestGPT (pentestgpt, ) and Cybench (zhang2024cybench, ) use a human in the loop, both to give suggestions to the LLM and to execute the commands the LLMs output. PentestGPT, uses state-of-the-art prompting strategies to improve the ability for LLMs to help humans in solving CTF challenges. PentestGPT outputs suggestions for high-level tasks (e.g., try to find vulnerabilities on the web server) and command line instructions (e.g., an nmap command). The human operator then manually executes the suggested action and reports back the result. Cybench explores how humans can specify low-level sub-tasks (e.g., what file contains the vulnerability) to improve the efficacy of LLMs outputting the correct bash command.
In summary, existing work, both in autonomous and human assisted LLMs, have shown preliminary promise for small CTF style security challenges. However, our understanding of how LLMs perform at executing multistage network attacks is limited. [^1] Some efforts have anecdotally claimed that LLMs may not be capable of multistage network attacks: “In tests of autonomous cybersecurity operations Llama 3 405B showed limited progress in our autonomous hacking challenge, failing to demonstrate substantial capabilities in strategic planning and reasoning over scripted automation approaches (cyberseceval3, ).”
2.2. Multi-stage Attack Evaluation Methodology
We implement 10 multistage attack environments ranging from 25–50 hosts with brief overviews in Table 1 and detailed descriptions in Appendix A. The environments are inspired from a mix of public reports of real-world attacks (equifax_report, ; colonial_pipeline_techtarget, ), common topologies (ciscoEnterpriseNetwork, ; ibmTreeNetwork, ), or used in prior work (mirage, ; ringNetwork2013technique, ; ferguson2021_deception_psychology, ; ciscoEnterpriseNetwork, ; dumbbellNetwork, ). The goals of these environments are either to exfiltrate critical data or access critical network hosts. Unlike CTF challenges, all of the environments are multistage, attackers have to conduct a variety of tasks to achieve the goal such as scanning networks, identifying vulnerabilities, exploiting remote services, escalating user privileges, and exfiltrating data. For instance, the Equifax-inspired environment requires attackers to execute over 246 unique tasks (we formally define a task as a sequence of commands in Sec. 3).
In order to systematically evaluate LLMs ability at conducting multistage attacks we require an LLM-agnostic tool that autonomously conducts attacks on a given environment. Since the autonomous tools are closed-source (cyberseceval3, ; o1systemcard, ; anthropic_AISI, ), we replicate prior work by building an LLM agnostic tool that follows the same process  [^2] This tool will also be open-sourced upon publication.. We instruct the LLM to attack the environment with the goal to exfiltrate any data and gain access to any critical hosts. In the prompt, we also include the external ip address range of the environment. Finally, we instruct the LLM to output specific command line tasks to execute. Then, our tool extracts the commands and executes them on a Kali host that has access to the environment. All the tools required to execute the multistage attacks are preinstalled on the Kali computer, and the computer has the top 10 most common attacker tools preinstalled.
We also validate that the state-of-the-art prompting strategies in PentestGPT (pentestgpt, ) do not realize end-to-end attacks. Since PentestGPT requires a human operator, we manually evaluate PentestGPT by inputting the goal prompt used for the autonomous system. Then we manually enter the commands into the attacker’s Kali host. To control for the human operator assisting the LLM, we only execute concrete commands from PentestGPT and if a concrete command is not given, we ask once to supply a concrete command. If no concrete command is given after asking, we end the trial.
We evaluate the autonomous command line attack tool across 3 LLMs across 10 different environments with 5 trials for each pairing. We also manually evaluate PentestGPT across all 10 environments with 3 trials each [^3] Due to manual effort and little variance between trials we use fewer trials.. We were unable to evaluate o1, a state-of-the-art “reasoning models”, because the public API has a safeguard that prevents o1 from executing attacks.
Figure 2. Across all LLMs and environments, LLMs were unable to fully realize an end-to-end multistage attack. Only Sonnet 3.5 was able to exfiltrate a single file in the 4-layer chain environment.
2.3. Findings
In Fig. 2, we measure LLM success by either not achieving a single goal (e.g., not exfiltrating a single file), partially succeeding at least once (e.g., exfiltrating at least one file), and fully succeeding (e.g., exfiltrating all files). We show this success metric for three state-of-the-art LLMs and PentestGPT for all 10 environments.
Across all evaluated LLMs and environments, we find that LLMs were unable to fully realize an end-to-end multistage attack. Only one LLM, Sonnet 3.5, was able to partially succeed in a single trial by exfiltrating 1 out of 25 files in the 4-Layer Chain environment. The human-assisted PentestGPT was unable to partially succeed in any environment; event state-of-the-art prompting strategies in PentestGPT do not help in this setting.
Takeaways: Prior work evaluates LLMs in single stage CTF challenges and showed preliminary promise. However, it is unclear how LLMs perform at more realistic multistage network attacks. We find that across several multistage environments and LLMs, LLMs are not able to realize end-to-end attacks. Additionally, we show that the state-of-the-art prompting strategies do not improve the ability for LLMs to conduct multistage attacks in these environments.
3. Why do LLMs struggle with multistage attacks
A natural question, then, is why did the LLMs fail to execute these multistage attacks. Unfortunately, prior work only offers very high-level guidance in this regard; e.g.,
  • Although the models often have good insights, they sometimes execute on them poorly. The agents also sometimes fail to pivot to a different strategy if their initial strategy was unsuccessful, or they miss a key insight necessary to solving the task (o1systemcard, ).
To shed light on these failure modes and inform our design, we use a first-principles approach using attack graphs (attack_graph_og, ; OG_MULVAL, ). At a high level, a multistage attack entails a complex end goal, where an attacker needs to break down the complex goal into a number of intermediate states. This is precisely what the attack graph formalism offers, as it provides a formal foundation for modeling attacker end goals, sub goals, intermediate states, and candidate actions to achieve intermediate states (attack_graph_og, ; OG_MULVAL, ). As we will see next, using the attack graph formalism helps us shed light on when and how LLMs failed in the multistage attack.
Figure 3. Part of the Equifax-inspired environment’s attack graph. The Equifax attack graph is complex and has 246 unique states.
3.1. Preliminaries
Formally, an attack graph is defined as $G=(S,A,S_{o},S_{g})$ where $S$ is a set of states, $A\subseteq S\times S$ is the set of actions (directed edges) representing transitions between these states, $S_{g}\subseteq S$ is the set of goal states, and $S_{o}\subseteq S$ is the set of initial states (attack_graph_og, ). Intuitively, in the attack graph, the nodes are attacker states (e.g., gained access to web erver) and the edges are attack actions (e.g., exfiltrate data). Attack graphs can be complex and consist of many intermediate states. Fig. 3 shows an example attack graph of the Equifax-inspired environment, which has 246 unique states. [^4] In this case, there are 48 goal states corresponding to the exfiltrating data from each database.
With an attack-graph formulation we can systematically analyze why the LLM attacker failed. We can measure when an attacker failed by identifying the states the attacker was able to achieve (e.g., found external hosts) and which ones they failed to achieve (e.g., unable to gain initial access). And, we can identify how an attacker failed by analyzing the edges in the attack graph. For instance, an LLM may have been unable execute an action (i.e., an edge) because a command had the incorrect parameters.
To execute these analyses, we need to incorporate the concept of a command into the attack graph. Each action $a\in A$ is composed of a sequence of commands. A single command is defined as a function $c:(h,n,p)\mapsto o$ where $h$ is the host on which the command is run, $n$ is the name of the command, $p$ are the parameters of the command, and $o$ is the output of the command.
Each action $a$ is a finite sequence of commands: $a=(c_{1},c_{2},\dots,c_{k})$ where each $c_{i}:(h_{i},f_{i},p_{i})\mapsto o_{i}$. We define a successful attack path, where an attacker achieves all of their goals, as $\pi=(s_{0},s_{1},\dots,s_{n})$ such that $S_{g}\subseteq\{s_{0},s_{1},\dots,s_{n}\}$.
With these preliminaries, next we describe how we use the attack graph formulation to analyze why LLMs failed in multistage attack.
3.2. Mapping LLM commands to attack graphs
Our first step is to logically map the LLM actions to the corresponding environment’s attack graph. Given that the command logs for a single trial is often thousands of lines long, manual analysis is intractable. To this end, we develop a framework to heuristically map LLM commands to ideal attack graphs. [^5] We do not claim these heuristics enable perfect mappings; these are sufficient to automatically analyze our logs and provide key insights.
Figure 4. Percentage of attack graph states achieved by LLMs using shell commands. Across all environments, LLMs are only able to achieve 1–30% of the states in the attack graph.
First, for each trial, we identify the states in the attack graph the LLM achieves. We identify these states by searching for keywords in command outputs. For instance, we search for relevant IP addresses to identify the number of hosts discovered and relevant CVEs to identify the number of vulnerabilities found. In Fig. 4, we show the maximum percentage of attack states achieved by each LLM in all 10 environments. Across all environments, LLMs are only able to achieve 1–30% of the states in the attack graph.
Now, we explore when and why LLMs are unable to achieve states in the attack graph. To this end, we use two environments as illustrative examples where there exists only a single successful attack path ($\pi$): the Equifax-inspired and 4-Layer Chain environments 666 In the Equifax-inspired environment there are two correct attack paths, but the LLMs in Sec. 2 never achieve the state where the paths diverge. . For each of these environments, we manually create an implementation of a successful attack using a sequence of commands $C_{\text{man}}=(c_{1},c_{2},\dots,c_{m})$.
Figure 5. We categorize LLM commands into 3 categories: irrelevant commands ($C_{I}$), relevant commands with correct implementations ($C_{\text{RC}}$), and relevant commands with incorrect implementations ($C_{\text{RI}}$). The two failure modes of LLMs are the commands in $C_{I}$ and $C_{\text{RI}}$.
Let $C_{\text{LLM}}=(c_{1},c_{2},\dots,c_{m})$ be the sequence of commands generated by the LLM. As seen in Fig. 5, we check each command in $C_{\text{LLM}}$ for two failure modes: (1) irrelevant commands $C_{I}$ and (2) relevant but incorrectly implemented commands $C_{\text{RI}}$.
The first possible failure mode of LLMs is irrelevant commands $C_{I}$. For each command $c_{j}\in C_{\text{LLM}}$, we call the command irrelevant if the command’s host, $h_{j}$, and command’s name, $n_{j}$, do not appear as part of any command in $C_{\text{man}}$. For example, LLMs sometimes output a hydra command, a tool for brute forcing passwords, but brute forcing passwords is not possible in either environment so we consider the command irrelevant. We manually inspect commands in $C_{I}$ to check if they correspond to an alternate implementation that we did not consider. In the two environments we analyze, no such alternate implementations exist and as a result all of the commands in $C_{I}$ are not part of actions in the attack graph.
The other failure mode involves commands that are relevant but have incorrectly implemented commands $C_{\text{RI}}$. We identify these commands by considering the set of relevant commands $C_{R}=C_{\text{LLM}}\setminus C_{I}$. In the set of relevant commands, we notice that several commands could have lead to useful attacker states in $S$, but did not because they were implemented incorrectly. For instance, it was common for LLMs to output nmap commands that could not identify vulnerabilities due to incorrect command parameters.
From the relevant set $C_{R}$, we identify those commands that correctly implement the intended action. Specifically, a command is tagged as correctly implemented if both of the following hold:
  1. (1)
    The parameters $p_{i}$ contain the correct parameters required by the implementation. For example, if the realization requires an nmap scan with a specific parameter set, the command must include these exact parameters.
  2. (2)
    Even with the correct parameters, the command can be implemented incorrectly because non-key parameters or syntax mistakes can result in a wrong implementations (e.g., command output is an error). To filter these commands out, we verify the output $o_{i}$ contains essential keywords or patterns that match the known correct output of the intended step. For instance, detecting a specific CVE identifier in the output indicates that the step was correctly executed.
The commands that satisfy the above criteria form the set of relevant and correctly implemented commands $C_{\text{RC}}$. All other relevant commands are considered incorrect implementation commands $C_{\text{RI}}=C_{R}\setminus C_{\text{RC}}$. Similar to $C_{\text{RI}}$, none of the commands in $C_{\text{RI}}$ are part of actions in the attack graph. The set of relevant but incorrect implementation commands $C_{\text{RI}}$ is this second failure mode.
Figure 6. In the Equifax-inspired and chain environments, 28–81% of the commands are irrelevant to the attack graph ($C_{I}$). Furthermore, 9–32% of the commands are are implemented incorrectly ($C_{\text{RI}}$).
3.3. Analysis results
We map all LLM outputs for all five trials for two environments. We only map LLM outputs for two environments because of the large manual effort in creating $C_{\text{man}}$. We choose the Equifax-inspired environment and the 4-Layer Chain environment because these are the environments the LLMs performed the best and worst in.
We show the percentage of LLM commands in each category in Fig. 6: relevant commands (i.e., successful commands that led to states in the attack graph), irrelevant commands (i.e., the first failure mode), and relevant commands with incorrect implementations (i.e., the second failure mode).
Commands irrelevant to attack (Failure mode 1): Across the LLMs and environments, 28–81% of commands are irrelevant to the multistage attack ($C_{I}$) shown in Fig. 6.
First, LLM commands failed to achieve attack graph states because many were irrelevant to the multistage attack. For instance, the LLMs tried brute forcing SSH credentials, finding misconfigured files, or exploiting non-exploitable services.
Incorrect command implementations (Failure mode 2): Across the LLMs and environments, 9–32% of the relevant bash commands are are implemented incorrectly ($C_{\text{RI}}$) shown in Fig. 6.
LLMs also struggled to correctly implement shell commands. For instance, LLMs try to find vulnerabilities using nmap. But the command had incorrect parameters and the scan reports that there were no vulnerabilities. In a multistage attacks these errors are critical as they cause cascading failures.
Furthermore, of the 235 relevant and correct implementation commands, 229 of them were nmap scans that identified hosts (not vulnerabilities). This suggests that LLMs are able to identify hosts in networks but struggle to identify and execute other types of commands required for multistage attacks.
Takeaways: Using the attack graph formalism sheds light on two critical failure modes of LLMs in multistage attacks: incorrect implementation of commands and executing commands irrelevant to the attack. These are critical failure modes that can have cascading impacts in the context of a complex multistage attack.
4. Design of Incalmo
To address the failure modes in Sec. 3, we present Incalmo, a high-level abstraction layer that sits between the LLM and the environment to be tested, shown in Fig. 7. Rather than the LLM directly outputting low-level shell commands (e.g., nmap scans, Metasploit exploits), we have LLMs output high-level tasks and queries (e.g., infect a host, scan a network, find a path to a host). Incalmo translates these tasks and queries into their corresponding low-level primitives, executes the primitives, appends the results to the prompt, and requests another task or query from the LLM.
Figure 7. Rather than LLMs directly outputting low-level shell commands, LLMs interact with Incalmo with high-level tasks or queries. Incalmo translates the tasks and queries into low-level primitives and returns the result.
4.1. The case for more abstraction
In some sense, seeing LLMs struggle to reason about multistage network attacks is not surprising—LLMs continue to struggle with complex reasoning for many other domains too (gao2023pal, ). Other domains have partially addressed these problems by offloading the solution step to high-level frameworks and APIs (gao2023pal, ; lu2024chameleon, ; schick2024toolformer, ). For instance, to answer a user’s question, LLMs sometimes use Wikipedia’s high-level API instead of generating low-level web requests (schick2024toolformer, ). Or in another case, LLMs use Bing’s high-level API to search the web for relevant information  (lu2024chameleon, ).
We observe a similar parallel in our problem setting. We argue the need for a high-level abstraction layer that the LLM can use to offload the solution and reasoning steps for multistage network networks. However, to the best of our knowledge, no such framework or high-level API exists in the security domain. Indeed, as we saw, prior work uses LLM as-is or add human-in-the-loop reasoning, and this is fundamentally limited.
To this end, we introduce Incalmo, a high-level abstraction layer for multistage attacks. Before we describe the key components of Incalmo and how it addresses these failure modes, we outline system requirements.
4.2. Requirements
We outline three key requirements that any multistage abstraction layer should satisfy:
@@ 4.2. Requirements
    LLM agnostic: New types of LLMs are constantly being created. We want our abstraction layer to work with any type of LLM. For example, we want to quickly evaluate and compare LLMs of different sizes.
@@
    Environment agnostic: Real networks often differ in topology, vulnerabilities, and configurations. We want an abstraction layer that remains independent of the specific environment.
@@
    Extensible: As new attack capabilities appear, the abstraction layer should be extensible for new techniques.
4.3. Detailed Design
The design of Incalmo builds on two key ideas. Our first insight is that we can use the attack graph formalism to avoid executing commands irrelevant to a multistage attack (3.3). Our second insight is to raise the level of abstraction for the commands that LLMs output to minimize the chances of incorrect implementations (3.3). That is, we instruct the LLM to output high-level tasks that Incalmo translates to correct low-level commands, instead of the LLM writing error-prone shell commands which results in dead ends in a multistage attack.
LLMs interact with Incalmo with two interfaces: a task or a query. A task is a high-level action to execute on hosts the attacker has access to (e.g., exfiltrate data from a host). A query is a request for information from Incalmo (e.g., find an attack path to a host).
To execute tasks and queries, Incalmo has three key modules:
@@ 4.3. Detailed Design
    An action planner to assist LLMs with the implementation of commands (3.3). The action planner translates high-level tasks into low-level primitives. Additionally, the action planner is extensible to support new attacker capabilities.
@@
    An attack graph service to assist LLMs in outputting relevant commands (3.3). LLMs and Incalmo’s modules can query the service to identify potentially relevant tasks to execute.

### v3

@@ II-A Related work on cyber-offense systems
Dumbbell B Has the same topology as Dumbbell A. Each web server’s credentials and the data on each database requires privileged access. Exfiltrate data from 15 databases. 30
II-B MHBench: A multi-host attack evaluation benchmark
To address the gap in evaluating LLM-based systems on multi-host attack challenges, we design and implement an open-source benchmark MHBench. MHBench has 10 multi-host attack environments, with summaries in Table II and detailed descriptions in Appendix A. The attacker goals are either to exfiltrate critical data or access critical network hosts.
We design MHBench to be diverse along key dimensions: (1) network size and topology, (2) type of vulnerabilities, and (3) attack complexity. For (1), MHBench currently includes many small enterprise environments ranging from 25 to 50 hosts. We use diverse network topologies such as tree-like topologies [2, 3], topologies from prior work such as “Star”, “Chain”, and “Dumbbell” [32, 58, 18, 2, 34], and topologies from public reports of real-world attacks [37, 29].
For (2), MHBench includes diverse vulnerabilities such as common misconfigurations (e.g., plain text credentials), remote code execution vulnerabilities (e.g., Apache Struts CVE-2017-5638), and privilege escalation vulnerabilities (e.g., vulnerable sudo CVE-2021-3156). Several of these vulnerabilities have been used in real-world attacks [20, 29].
Finally, the MHBench’s environments vary in terms of attack complexity w.r.t the number of required steps for a successful attack. The environments range from 125 to 246 unique attack graph states. We formally define this in Sec. III. To put this in context, the challenges used in PentestGPT’s evaluation require fewer than 10 states. We also vary the number of subgoals the attacker has to tackle to ensure full success. For instance, the Enterprise B environment requires exfiltrating data from 9 database instances while the Equifax-inspired environment requires exfiltrating data from 48 database instances.
Benchmark metrics. We evaluate LLM-based attackers on the following metrics:
@@ II-B MHBench: A multi-host attack evaluation benchmark
    Attack success: We consider a full success as an LLM reaching all goal states (e.g., exfiltrate all 48 data files in the Equifax-inspired environment) and no success as an LLM failing to reach any goal states. For partial success, we envision two metrics. First, we quantify the partial number of success goals (e.g., exfiltrated 12 out of 48 data files). Second, we create a reference attack graphs [53] for each environment. We use the graphs to compute a coverage metric of the percentage of states reached along the reference attack path(s).
@@
    Time: We capture the total attack execution time and the time taken to reach each goal state.
@@
    Cost: We measure the total cost in terms of the number of input and output tokens used by the LLM.
Limitations. MHBench only contains known vulnerabilities and does not include 0-day vulnerabilities. In addition, MHBench does not have Windows machines. We do note that MHBench is extensible. We will add new environments and make it easy for the community to contribute new challenges.
II-C Evaluating existing LLM-based approaches on MHBench
Next, we evaluate state-of-art LLM-based systems at executing multi-host attacks using MHBench. For this section, we only focus on the success metrics. We consider two categories of existing LLM-based cyber-offense approaches as baselines:
@@ II-C Evaluating existing LLM-based approaches on MHBench
    Autonomous shell-based systems: The first baselines we consider are shell-based systems, systems that instruct the LLM to output shell commands to execute. [43, 59, 67, 60]. Then, a program extracts the commands and executes them on a Kali Linux host that has access to the environment, and equipped with the required tools to execute the attacks. These systems use state-of-the-art “reasoning” techniques such as chain-of-thought [60, 61], ReAct [65], and self-reflection [60, 13]. We study using two specific type of prompts: CyberSecEval3’s [60] prompt and Domain-shell, a prompt we create in collaboration with a domain expert at a leading LLM provider. Both prompts can be found in Appendix C. Domain-shell’s prompt is shorter and has less environment specific information than CyberSecEval3’s prompt. In both prompts, we include the external IP address range of the environment.
@@
    PentestGPT-like systems: We also evaluate state-of-the-art LLM-based systems that break complex security tasks into sub tasks (e.g., [27, 13, 67]). In particular, we evaluate PentestGPT [13], a tool with many of these prompting strategies (and all of the prior “reasoning” techniques). Since PentestGPT requires a human operator, we manually evaluate PentestGPT by telling it the external IP address range of the network. Then we manually enter the commands into the attacker’s Kali Linux host. To control for the human operator assisting the LLM, we only execute concrete commands from PentestGPT and if a concrete command is not given, we ask once to supply a concrete command. If no concrete command is given, we end the trial. (Even though this is technically not an autonomous system, we evaluate it to see if the techniques can help address the multi-host network scenario.)
For the shell-based system baselines, we consider three state-of-the-art (at the time of writing) LLMs: Sonnet 3.7 Thinking, GPT 4o, Gemini 2.5 Pro. [^2] We were unable to evaluate OpenAI’s “o” models because the public API has a safeguard that prevents them from executing attacks. Later in Sec. VII, we consider six other LLMs that perform worse on standard LLM benchmarks. All three models are evaluated for all 10 environments in MHBench with 5 independent trial runs. Since PentestGPT requires manual effort and recommends using GPT4o, we use it across all 10 environments with 3 trials each [^3] Due to manual effort and little variance between trials we use fewer trials.. To put these results in context, we also evaluate Caldera, a non-LLM tool, at executing multi-host attacks.
Figure 2: Across all LLMs and environments, LLMs were unable to fully realize an end-to-end multi-host attack. Only Sonnet 3.7 Thinking with the Domain-shell was able to partially succeed in the 4-layer chain environment.
Findings. Across all evaluated LLMs and environments, we find that LLMs were largely unable to realize multi-host attacks. Only one LLM, Sonnet 3.7 Thinking with the Domain-shell prompt, was able to partially succeed by exfiltrating 11 out of 25 files in the 4-Layer Chain environment, shown in Fig. 2. The human-assisted PentestGPT was unable to partially succeed in any environment; even state-of-the-art prompting strategies in PentestGPT do not help in this setting.
We have tested other models such as DeepSeek and Llamma 3 but do not show these for brevity444 In our experiments, these models do not follow instructions and are unable to execute shell commands in the correct format. . As models get released, we plan to update our benchmark “scorecard” (Fig. 2). However, we believe that incremental updates will not fundamentally change the results; prompt hacks are unlikely to fix the issues we highlight next.
Takeaways. We find that across several multi-host environments and LLMs, LLMs are unable to execute end-to-end attacks. Even state-of-the-art prompting strategies do not enable LLMs to conduct multi-host attacks in these environments.
III Understanding how LLMs fail in MHBench
The previous results showed that LLMs failed in the end-to-end success metric. A natural question, then, is how and why did the LLMs fail. Unfortunately, existing efforts in using LLMs for cyber offense do not offer actionable insights in this regard [43]. In this section, we use an attack graph formalism to define and understand these failure modes [53, 45]. This helps shed light on how the LLMs failed and also informs our design choices in the next section.
Figure 3: The figure shows the Equifax attack graph is complex and has 246 unique states.
III-A Preliminaries and Methodology
A multi-host network attack is complex and involves multiple intermediate subgoals, strategic planning, and coordinated actions at each step toward the final target(s). Attack graphs give us a formal foundation for modeling attacker end goals, sub goals, intermediate states, and candidate actions [53, 45].
Formally, an attack graph is defined as $G=(S,A,S_{o},S_{g})$ where $S$ is a set of states, $A\subseteq S\times S$ is the set of actions (directed edges) representing transitions between these states, $S_{g}\subseteq S$ is the set of goal states, and $S_{o}\subseteq S$ is the set of initial states [53]. Intuitively, the nodes are attacker states (e.g., gained access to web server) and the edges are attack actions (e.g., exfiltrate data). We define a successful attack path, where an attacker reaches all of their goals, as $\pi=(s_{0},s_{1},\dots,s_{n})$ such that $S_{g}\subseteq\{s_{0},s_{1},\dots,s_{n}\}$. Fig. 3 shows a simplified attack graph for the Equifax-inspired environment, which has 246 unique states. [^5] In this case, there are 48 intermediate goal states corresponding to the exfiltrating data from each database instances.
For each environment, we manually create a reference attack graph and a minimal sequence of commands for an attack: $C_{\text{man}}=(c_{1},c_{2},\dots,c_{m})$.666 For most of the environments, there is only one successful attack path. Here, a single command is defined as a function $c:(h,n,p)\mapsto o$ where $h$ is the host on which the command is run, $n$ is the name of the command, $p$ are the parameters of the command, and $o$ is the output of the command. Using $C_{\text{man}}$, we can analyze when and how the LLM attacker failed. First, we can identify which “nodes” (states) the attacker was able to reach (e.g., found an external host) and which ones they failed to reach (e.g., unable to gain initial access). Second, by analyzing attacker actions at each “edge”, we can check if this edge successfully induced a transition to the next state or identify how it failed.
III-B Analysis results
First, we show that LLMs did not even reach relevant states. Then, we analyze the commands to understand how, and why, they failed to reach the relevant states.
Figure 4: Percentage of attack graph states reached by LLMs using shell commands. Across all environments, LLMs are only able to reach 1–30% of the states in the attack graph.
Attack graph states reached. To track if the LLMs reached some of the relevant attack states, we use the following heuristic. For each state $s$ in the attack graph, we use $C_{\text{man}}$ to track the relevant output $o_{\mathit{s}}$ triggered when the state $s$ was reached. For example, a correct implementation of a vulnerability scan will output a specific CVE, denoting we correctly discovered the CVE. Given a sequence of commands generated by the LLM $C_{\text{LLM}}=(c_{1},c_{2},\dots,c_{m})$, we can match the command outputs against the reference $o_{\mathit{s}}$ values to check if the state $s$ was reached. That is, for each command $c_{j}\in C_{\text{LLM}}$, we check if $o_{j}$ matches the keywords associated with $o_{\mathit{s}}$. If it matches, we consider the LLM to have reached this state $s$ [^7] We acknowledge that there could be alternative ways to achieve a state that do not contain these keywords. We do our best effort to manually review the logs to ensure this is not the case. Across all environments, LLMs are only able to reach 1–30% of the states as seen in Fig. 4. The figure shows the maximum (across trials) of the percentage of attack states reached by LLMs in each MHBench environment.
Figure 5: In the Equifax-inspired and chain environments, 47–90% of the commands are irrelevant to the attack graph ($C_{I}$). Furthermore, 6–41% of the commands are are implemented incorrectly ($C_{\text{RI}}$).
Command analysis. The above result suggests that the LLMs did not make substantial progress in the attack graph. To further understand how they failed, we analyze the commands in $C_{\text{LLM}}$. For this analysis, we focus on two exemplar environments: the Equifax-inspired and 4-Layer Chain environments.
We identify two modes of failures: